On the Standford-Binet test, the mean IQ is 100. A class of 21 kindergarten pupils were tested with a resulting mean of 100 and a standard deviation of 8.06. What is the standard error of the 90% confidence interval of the average IQ of the kindergarten pupils?



1
Expert's answer
2022-02-28T11:01:24-0500

"SE=\frac{s}{\sqrt{n}}=\frac{8.06}{\sqrt{21}}=1.76."